### v3

# Alt-MoE:A Scalable Framework for Bidirectional Multimodal Alignment and Efficient Knowledge Integration

Abstract
Multimodal learning has advanced significantly by aligning different modalities within shared latent spaces, enabling tasks such as cross modal understanding and generation. Current alignment strategies in multimodal learning primarily include direct alignment using pre-trained or unified encoders and single-directional alignment via modality-specific connectors. Direct alignment struggles to fully leverage rich intra-modal knowledge, often requiring extensive training data to achieve cross-modal representation. Meanwhile, single-directional alignment methods, despite leveraging pre-trained knowledge, restrict task adaptability and hinder the model’s ability to capture bidirectional relationships, leading to incomplete knowledge fusion and underutilization of complementary modality-specific information. To address these limitations, we introduce Alt-MoE, a scalable multimodal alignment framework that employs a mixture of experts (MoE) model as a multi-directional connector across modalities. By utilizing a sequential alternating one-way alignment strategy, Alt-MoE iteratively refines the model to achieve bidirectional alignment. Alt-MoE operates in latent space, enabling efficient vector pre-storage and real-time retrieval via MoE, optimizing large-scale data processing. Extensive empirical studies demonstrate that Alt-MoE achieves competitive performance on cross-modal retrieval and visual question answering by integrating diverse modality-specific knowledge, generalizing to unseen data, and easily scaling to new tasks and modalities through dynamic adjustment of MoE capacity and expert activation.
1 Introduction
Figure 1: Primary strategies for aligning modalities: Figure (a). Direct Alignment Framework: This framework illustrates two types of alignment. On the left, a dual-encoder approach is shown, where an x-encoder and a y-encoder independently process inputs $x$ and $y$, producing representations $s_{x}$ and $s_{y}$. A similarity function $D(s_{x},s_{y})$ is used to align the two modalities directly. On the right, a unified multimodal encoder jointly processes $x$ and $y$, producing a fused representation $s_{xy}$. The training objective combines two loss functions: $L_{itm}$, focusing on image-text matching, and $L_{lm}$, which corresponds to the language modeling loss. Figure (b). Unidirectional Alignment Framework: In this framework, an x-encoder processes input $x$ to produce a representation $s_{x}$, which is then refined by a predictor. $I$ provides additional information to guide the predictor in aligning refined representations with a large language model (LLM). The LLM is trained with the language modeling loss $L_{lm}$, using input $y$ as the target. This framework focuses on aligning the modality $x$ unidirectionally to the language modality through the LLM. Figure (c.) Alt-MoE models cross-modal bidirectional alignment by alternately executing unidirectional alignment tasks in the latent space (detail in 4.3). $s_{x}$, $s_{y}$ represent the encoding results of the $x$-encoder and $y$-encoder, respectively. $t$, $t+1$ represent the time steps. $L$ denotes the loss function (prediction and contrastive loss). $z_{x2y}$ and $z_{y2x}$ represent the cross embeddings (detail in 4.3).
Achieving human-level intelligence, a central goal of Artificial General Intelligence (AGI), requires building hierarchical world models capable of abstract reasoning. These models enable agents to predict, plan, and reason about complex, multi-layered environments. Abstract reasoning, particularly in high-dimensional latent spaces, is crucial for cross-modal understanding and generation. It requires integrating diverse sources of information across multiple modalities to construct unified, high-level representations. These representations must transcend individual modalities, enabling models to capture complex, abstract relationships [22, 8].
However, current multimodal learning often fall short in effectively capturing and utilizing this abstract reasoning across modalities. Current multimodal learning frameworks often rely on two primary strategies for aligning modalities: direct alignment using pre-trained or unified multimodal encoders and single-directional alignment through modality-specific connectors. Direct alignment methods, such as CLIP [37] and ImageBind [13], involve aligning modalities by mapping them into a shared space using pre-trained encoders. VLMo [4] and ONE-PEACE [40] utilize a unified multimodal encoder to integrate and model interactions across different modalities. On the other hand, single-directional alignment strategies, like BLIP-2 [25] and LLaVA [33], focus on connecting specific modalities through specialized components or connectors, often operating in a one-way alignment direction. These approaches aim to establish coherent cross-modal representations, but each comes with its limitations in capturing the full complexity of multimodal relationships.
Direct alignment of multiple encoders excel in generalization, and efficient cross-modal retrieval [25, 42, 13, 47]. Unified multimodal encoder facilitates deeper interactions across modalities and is particularly advantageous in scenarios requiring joint reasoning and holistic understanding [41, 40, 4]. Nevertheless, both approaches share limitations: they often miss fine-grained intra-modal details, struggle with tasks that demand abstract reasoning, and rely heavily on large-scale pre-training data. In text-guided visual representation learning, captions offer an abstract view of images, often missing pixel-level details crucial for complex visual reasoning. This limitation challenges direct alignment methods in retaining fine-grained intra-modal features, restricting their deep visual understanding capabilities [35].
Another method employs single-directional alignment through modality-specific connectors, such as LLaVA [33] and BLIP-2 [26]. These methods connect different modalities via specialized components, aligning them in a single direction. While this approach facilitates efficient cross-modal interaction and preserves unimodal richness, it can limit the model’s ability to capture complex, bidirectional relationships necessary for abstract reasoning. As a result, while it excels in specific tasks, it may struggle with more intricate, abstract reasoning that requires deeper intermodal dependencies.
Building upon these observations, we introduce Alt-MoE, a novel multi-modal alignment framework. Alt-MoE employs a unified, multi-directional MoE as cross-modal connector, enabling any-to-any modality alignment. This framework integrate the rich intra-modal knowledge of pre-trained models, preventing the loss of fine-grained information while avoiding high computational costs and data requirements. By employing Alternating Gradient Descent (AGD), bidirectional alignment decomposes into alternating unidirectional tasks, ultimately achieving bidirectional alignment. This method not only mitigates the risks of incomplete information and abstract reasoning deficiencies inherent in single-directional alignment but also allows the model to incorporate multiple loss functions, achieving more robust cross-modal alignment. Alt-MoE introduces a paradigm shift by exploring latent dependencies, enabling more dynamic and abstract forms of reasoning.
Further expanding on this, Alt-MoE is a new multi-modal paradigm which studies the dependency of the unobserved part ($y$) on the observed part ($x$) in their latent space. This enables the model to explore different hypothetical scenarios, plan actions, or make inferences about unobserved parts of the world based on observed data, thus facilitating more sophisticated and generalized reasoning. Experts within the MoE can operate at different levels of abstraction, ensuring that each modality is processed in a way that suits its nature while maintaining coherence in the overall model.
To assess the model’s performance, we achieved superior results across multiple tasks (e.g., VQA, image-text retrieval, audio-text retrieval, image classification) and datasets. Additionally, we validated the model’s generalization on unseen data by testing its performance on zero-shot audio-text retrieval. Experimental results demonstrate that our Alt-MoE easily achieves modality and task scaling while attaining competitive results compared to current state-of-the-art (SoTA) multi-modal studies. Furthermore, our approach is able to conduct large-scale on-line multi-modal retrieval tasks due to our unique architecture. We summarize our contributions as follows:
  1. 1.
    Alt-MoE introduces a modality-agnostic alignment method that reduces training and data costs while leveraging the rich knowledge in pre-trained models. This approach offers a new framework for multi-modal research, potentially enhancing efficiency and scalability across diverse modalities.
  2. 2.
    Alt-MoE leverages multi-directional MoE for cross-modal alignment in latent space, building a hierarchical world model that abstracts complex dependencies.
2 Related Works
2.1 End-to-end Multi-modal Learning
Recently, end-to-end multi-modal models employing various architectures have achieved outstanding performance. These architectures can be broadly categorized into several main types: Dual-encoder [37, 18]. unified-encoder [27, 18, 6]. Combining dual encoder and fusion encoder architectures integrate specialized layers into multi-modal models to enable deep cross-modal interactions [24, 25].
The majority of multi-modal learning method employ large-scale multi-modal datasets for end-to-end pre-training. However, as model scale continues to increase, several potential challenges emerge:d Firstly, the pre-training process may incur prohibitively high computational costs. Secondly, these models often struggle to adapt to novel modalities or tasks without extensive retraining. Moreover, most multi-modal approaches employ text-guided learning of visual concepts, which can limit the acquisition of fine-grained information inherent to each modality. A logical approach is to leverage existing pre-trained uni-modal foundation models [26, 35, 45].
2.2 Multi-modal Learning with Uni-modal Models
Recent trends in multi-modal learning have increasingly focused on integrating high-performance uni-modal models to achieve multi-modal alignment. Flamingo [2] integrates visual information into each layer of a frozen Large LLM through the use of cross-attention. BLIP-2 [26] introduces an additional vision-to-language adaptation module, Q-former, and proposes a two-stage training process to mitigate the challenges associated with learning vision-language alignment. However, both methods require substantial parameters and multi-modal data for cross-modal alignment.
Recent studies indicate a convergence of representations across modalities, providing evidence for the feasibility of developing advanced multi-modal models by connecting high-performance uni-modal models with lightweight parameters and data-efficient techniques [16]. For instance, LLaVA [33] achieved state-of-the-art performance by employing a two-layer multilayer perceptron (MLP). Similar architectures have subsequently proliferated across various domains [23, 45, 30, 29]. Alt-MoE further advances this concept by interconnecting diverse high-performance uni-modal models through a shared multi-directional MoE. We conducted extensive experiments focused on alignment efficacy, demonstrating the effectiveness of joint training across multiple modalities and directions.
2.3 Multi-modal Learning with MoE and AGD
Prior studies have investigated AGD-based multi-modal multi-task alternating training, revealing that the integration of diverse modalities, tasks, and resolutions can yield mutual benefits, thereby effectively enhancing the model’s generalization capabilities and cross-domain performance [1, 28]. we further extend this approach to integrate existing pre-trained high-performance uni-modal models, achieving overall alignment through alternating bidirectional alignment.
MoE-LLaVA [30] proposes MoE-Tuning, a strategy for Large Vision-Language Models that creates a sparse model with constant computational cost. While both MoE-LLaVA and Alt-MoE employ sparse MoE to connect high-performance uni-modal large models, Alt-MoE not only scales this approach to large audio-visual-language models but also achieves pairwise bidirectional alignment across modalities.
3 Problem Formulation
In this section, we first introduce the overall optimization objective of bidirectional alignment, followed by the optimization objective of unidirectional alignment, and finally the alignment loss and theoretical explanation. To clearly articulate the design rationale, we illustrate this with the example of image-text alignment.
  1. 1.
    Multi-modal alignment objective: Alt-MoE decomposes multi-modal bidirectional alignment into multiple unidirectional alignment subtasks, where MoE serves as a cross modal knowledge representation and fusion module and achieve multi-modal alignment by alternately aligning from image to text and text to image.
  2. 2.
    Alternating unidirectional alignment: Alt-MoE maximizes mutual information and minimizes conditional entropy to obtain independent and shared information for unidirectional alignment. Specifically, Alt-MoE adds embeddings as prior information to the input representations to guide the MoE in performing different pre-training tasks.
  3. 3.
    Information Decomposition and Alignment: we provide an information-theoretic explanation to elucidate the rationale behind the design of the training objective.
3.1 Multi-modal Alignment objective:
In this section, we will introduce the decomposition of the multi-modal bidirectional alignment objective. multi-modal alignment aims to align diverse modalities in a latent space by finding optimal parameters $\theta$ that minimize an alignment loss $\mathcal{L}_{align}$. This can be formulated as:
$\theta^{*}=\arg\min_{\theta}\mathcal{L}_{\text{align}}(\theta).$ (1)
By combining MoE $f_{\theta}(\cdot)$ parameterized by $\theta$ and AGD, which can decompose the optimization objectives for multi-modal alignment $\mathcal{L}_{align}$ into multiple unidirectional alignment subtasks, and then alternately execute each subtask at various time step to achieve overall alignment. Specifically, the image-text alignment can be decomposed as follows in the following Equation 2:
$\mathcal{L}_{i-t}(\theta)$ $=\mathcal{L}_{i\rightarrow t}(\theta_{i\rightarrow t})\oplus\mathcal{L}_{t\rightarrow i}(\theta_{t\rightarrow i}),$ (2)
$\theta$ $=\theta_{i\rightarrow t}\cup\theta_{t\rightarrow i},$
where $\mathcal{L}_{i\rightarrow t},\theta_{i\rightarrow t}$ represents the image-to-text ($t\rightarrow i$) alignment objective and parameter subset, $\mathcal{L}_{t\rightarrow i},\theta_{t\rightarrow i}$ represents the text-to-image ($i\rightarrow t$) alignment objective and parameter subset, $\oplus$ denotes an alternating optimization operation at various time step. Therefore, we decompose the image-text alignment into multiple unidirectional alignment optimizations and parameter subsets.
3.2 Alternating Unidirectional Alignment
In this section, we delineate the optimization objectives and parameter update procedures across various time steps $t$. By introducing AGD, we can alternately optimize unidirectional alignment at each time step $t$ with the goals of maximizing mutual information between image $I$ and text $T$, denoted as $I(I;T)$, and minimizing conditional entropies, denoted as $H(T|I)$ and $H(I|T)$. The details of the multi-directional MoE are provided in the supplementary materials.
Given time step $t$, the objective function is updated as Equation 3:
$\mathcal{L}_{i-t}=\begin{cases}{\mathcal{L}_{i\rightarrow t}}^{(t)}=-I(I;T)+H(T|I),&\text{if }t=2k\\ {\mathcal{L}_{t\rightarrow i}}^{(t)}=-I(I;T)+H(I|T),&\text{if }t=2k+1,\end{cases}$ (3)
where $k$ is a non-negative integer. Based on Equation 1 and 3, the overall optimization objective can be formulated as shown in Equation 4:
$\theta^{*}=\arg\min_{\theta}\left(-I(I;T)+\lambda\left(H(T|I)+H(I|T)\right)\right),$ (4)
where $\lambda$ is a weight parameter.
Given a set of parameters $\theta$, we alternate between image-to-text and text-to-image unidirectional alignment at different time steps. At different time steps, we update only a subset of the parameters: $\theta_{i\rightarrow t}$ for $i\rightarrow t$ alignment and $\theta_{t\rightarrow i}$ for $t\rightarrow i$ alignment. Ultimately, this process ensures that all parameters are updated, such that $\theta=\theta_{i\rightarrow t}\cup\theta_{t\rightarrow i}$. The overall update process can then be described by the following Equation 5:
$\theta^{t+1}=\begin{cases}\theta^{t}-\eta\nabla_{\theta_{i\rightarrow t}^{t}}\mathcal{L}_{i\rightarrow t}^{(t)},&\text{if }t=2k\\ \theta^{t}-\eta\nabla_{\theta_{t\rightarrow i}^{t}}\mathcal{L}_{t\rightarrow i}^{(t)},&\text{if }t=2k+1,\end{cases}$ (5)
where $\eta$ is the learning rate, $k$ is a non-negative integer.
3.3 Information Decomposition and Alignment
For image to text alignment, conditional entropy $H(T|I)$ and $H(I|T)$ represent modality-specific information in text and image respectively, measuring uncertainty in one modality after observing the other. Mutual information $I(T;I)$ quantifies shared information between image and text modalities, indicating how much knowing one reduces uncertainty about the other. For accurate image-text alignment, high mutual information (more shared content) and low conditional entropy (less modality-specific information) are desirable, ensuring strong semantic coupling between modalities.
Alt-MoE leverages the MoE router to automatically select different experts, optimizing for these two objectives. This approach helps decouple modality-specific information from shared information, potentially improving the balance between capturing unique modal features and cross-modal relationships.
Figure 2: Overview of Alt-MoE: Alt-MoE connects different modalities through multi-directional MoE and performs unidirectional alignment from image to text and text to image at different time steps, ultimately achieving overall bi-directional alignment. Here, we take image-text alignment as an example. The left part of the figure illustrates the process of image-text alignment, where we alternately perform image-to-text and text-to-image alignment on a unified multi-directional MoE. The specific structure of the multi-directional MoE is shown in the figure on the right. Different embeddings prompt the MoE to select different experts to perform contrastive learning or prediction. CL refers to Contrastive Learning, and Dual Pred denotes Bidirectional Prediction. To guide the MoE in performing different sub-tasks, we introduce modality embeddings (e.g., text embeddings, image embeddings). These embeddings are incorporated through cross-embedding mechanisms to effectively steer the tasks.
4 Methodology
In this section, we take image-text alignment as an example to provide a detailed introduction to the architecture of Alt-MoE. As shown in Figure 2, Alt-MoE is divided into three modules: the visual model (VM), the language model (LM), and the bidirectional alignment module MoE. We input the paired images and text $(I,T)$ into VM and LM respectively to obtain latent representations $z_{i},z_{t}$, and then perform multi-modal interaction in the fusion module.
4.1 Image and Text Encoding
Given a pair of image and text inputs $(I,T)$, we employ separate encoders to process each modality. The visual model (VM) $f_{v}(\cdot)$ encode the image $I$, while the language model (LM) $f_{l}(\cdot)$ encode the text $T$. This process results in latent representations $z_{i}$ and $z_{t}$ for the image and text, respectively. The encoding can be formally expressed as:
$z_{i}$ $=f_{v}(I),\quad z_{i}\in\mathbb{R}^{i},$ (6)
$z_{t}$ $=f_{l}(T),\quad z_{t}\in\mathbb{R}^{t},$
where $z_{i}\in\mathbb{R}^{i}$ is the image representation, and $z_{t}\in\mathbb{R}^{t}$ is the text representation.
4.2 Image and Text Encoding
Given a pair of image and text inputs $(I,T)$, we employ separate encoders to process each modality. The visual model (VM) $f_{v}(\cdot)$ encode the image $I$, while the language model (LM) $f_{l}(\cdot)$ encode the text $T$. This process results in latent representations $z_{i}$ and $z_{t}$ for the image and text, respectively. The encoding can be formally expressed as:
$z_{i}$ $=f_{v}(I),\quad z_{i}\in\mathbb{R}^{i},$ (7)
$z_{t}$ $=f_{l}(T),\quad z_{t}\in\mathbb{R}^{t},$
where $z_{i}\in\mathbb{R}^{i}$ is the image representation, and $z_{t}\in\mathbb{R}^{t}$ is the text representation.
4.3 Unidirectional Alignment
At different time steps $t$, Alt-MoE performs unidirectional alignment tasks using different parameter updates and optimization objectives according to Equations 3 and 5. To guide MoE in selecting different experts for different modalities and tasks, we set up trainable modality encodings $E_{T},E_{I}$ and trainable task encodings $E_{ce},E_{mi}$. Figure 3 illustrates the cross embedding process, where these embeddings are combined and added to $z_{t},z_{i}$ to guide the execution of different tasks.
[…]
Figure 3: Cross Embedding.
Specifically, we add modality embeddings to the input space $z_{i},z_{t}$ to enable the fusion module MoE to perceive the input modality, and then linearly project it into the d-dimensional common space $z^{c}$ as follows:
$z^{c}_{i}=W^{1}_{i}\cdot z_{i},\quad z^{c}_{t}=W^{1}_{t}\cdot z_{t},\quad z^{c}_{i},z^{c}_{t}\in\mathbb{R}^{d}$ (8)
where $W^{1}_{i}\in\mathbb{R}^{d\times i}$ and $W^{1}_{t}\in\mathbb{R}^{d\times t}$ are learnable projection matrices, $d_{i}$ and $d_{t}$ are the dimensions of the original image and text feature spaces respectively, and $d$ is the dimension of the common space.
In the common latent space $z^{c}$ with consistent dimensions, we update the loss function $\mathcal{L}_{i\rightarrow t},\mathcal{L}_{t\rightarrow i}$ at different time steps $t$ according to Equation 5. We add different task embeddings to $z^{c}$ to enable the MoE $f_{\theta}(\cdot)$ to perceive tasks. Then, at different time steps $t$, we perform prediction and contrastive learning between image and text to maximize mutual information and minimize conditional entropy. Here, we take image-text alignment as an example; the same applies to text-audio alignment. This process alternates between image-to-text and text-to-image tasks as follows:
For image-to-text (at even time steps $2k$):
$\mathcal{L}_{i\rightarrow t}=\mathcal{L}_{\text{CE}}(\hat{z}_{t},z_{t})+\mathcal{L}_{\text{MI}}(z^{\text{mi}}_{i},z_{t}),$ (9)
where:
$\hat{z}_{t}=W^{1}\cdot f^{1}_{\theta_{i\rightarrow t}}(z^{c}_{i}+E_{\text{ce}}),$
$z^{\text{mi}}_{i}=W^{2}\cdot f^{2}_{\theta_{i\rightarrow t}}(z^{c}_{i}+E_{\text{mi}}),$
$f_{\theta_{i\rightarrow t}}=f^{1}_{\theta_{i\rightarrow t}}\cup f^{2}_{\theta_{i\rightarrow t}},$
where $W^{1}\in\mathbb{R}^{d\times t}$, $W^{2}\in\mathbb{R}^{d\times t}$ is a learnable projection matrix that map the output to dimensions consistent with $z_{t}$.
For text-to-image (at odd time steps $12k+$ 1):
$\mathcal{L}_{t\rightarrow i}=\mathcal{L}_{\text{CE}}(\hat{z}_{i},z_{i})+\mathcal{L}_{\text{MI}}(z^{\text{mi}}_{t},z_{i}),$ (10)
where:
$\hat{z}_{i}=W^{2}_{t}\cdot f^{1}_{\theta_{t\rightarrow i}}(z^{c}_{t}+E_{\text{ce}}),$
$z^{\text{mi}}_{t}=W^{2}_{t}\cdot f^{2}_{\theta_{t\rightarrow i}}(z^{c}_{t}+E_{\text{mi}}),$
$f_{\theta_{t\rightarrow i}}=f^{1}_{\theta_{t\rightarrow i}}\cup f^{2}_{\theta_{t\rightarrow i}},$
where $W^{3}\in\mathbb{R}^{d\times i}$, $W^{4}\in\mathbb{R}^{d\times i}$ is a learnable projection matrix that map the output to dimensions consistent with $z_{i}$.
4.4 Loss Function
We adopt a unidirectional prediction approach at each time step to minimize conditional entropy. This method involves predicting either text features from image features or image features from text features, alternating between time steps. We define the prediction loss function using the $L2$ distance as follows:
$\mathcal{L}_{\text{CE}}=|\hat{z}_{t}-z_{t}|_{2}^{2}+|\hat{z}_{i}-z_{i}|_{2}^{2},$ (11)
where $|\cdot|_{2}^{2}$ denotes the squared $L2$ norm. $\hat{z}_{t}$ and $\hat{z}_{i}$ are the predicted text and image features, respectively, and $z_{t}$ and $z_{i}$ are the corresponding target features. At each time step, only one of these terms is active, depending on the prediction direction.
Here, we set $z^{mi}_{t}$ as $z^{T}$ and $z^{mi}_{i}$ as $z^{I}$and contrastive loss can be formulated as follows:
$\mathcal{L}_{\text{MI}}=\frac{1}{2N}\sum_{i=1}^{N}\Bigg{[}$ $-\log\frac{\exp(\text{sim}(z_{i}^{I},z_{i}^{T})/\tau)}{\sum_{j=1}^{N}\exp(\text{sim}(z_{i}^{I},z_{j}^{T})/\tau)}$ (12)
$-\log\frac{\exp(\text{sim}(z_{i}^{T},z_{i}^{I})/\tau)}{\sum_{j=1}^{N}\exp(\text{sim}(z_{i}^{T},z_{j}^{I})/\tau)}\Bigg{]}$
where:
  • $N$ is the number of image-text pairs in a batch.
  • $z_{i}^{I}$ and $z_{i}^{T}$ are the latent representations of the $i$-th image and its corresponding text, respectively.
  • $\text{sim}(z_{i}^{I},z_{j}^{T})$ is the cosine similarity between the latent representations $z_{i}^{I}$ and $z_{j}^{T}$.
  • $\tau$ is a temperature parameter that controls the sharpness of the similarity distribution.
5 Experiments and Results
We report Alt-MoE’s performance across various tasks, including image-text retrieval, speech-text retrieval, visual question answering, visual reasoning, and image classification on multiple datasets. Additionally, we validate Alt-MoE’s generalization capability on unseen data through zero-shot speech-text retrieval experiments. In the ablation study, we validated the impact of different Alt-MoE components through experiments on image-text retrieval tasks.
5.1 Datasets
The dataset details are provided in the supplementary materials. Image-text retrieval: COCO [31], Flickr30K [36]. Zero-shot audio-text retrieval: Clotho [11], Audiocaps [20]. VQA: VQA v2 [3], NLVR-2 [38]. Image Classification: ImageNet 1k [9].
5.2 Image-text Retrieval
Alt-MoE connects two uni-modal models through a multi-directional MoE, we train the multi-directional MoE on the COCO and Flickr30K, and then test it on the test sets. Table 1 shows the results of Alt-MoE on COCO and Flickr30K by integrating different LLMs such as LLAMA3-8b and Qwen2-7B with the LVMs Dinov2-Large. The results indicate that by integrating existing high-performance uni-modal models, Alt-MoE achieves superior performance.
Furthermore, by comparing different architectures of multi-modal models, Alt-MoE demonstrates superior data efficiency and parameter efficiency. Specifically, Alt-MoE has only 140M trainable parameters, which is significantly smaller than BLIP-2’s 1.2B trainable parameters.
The outstanding retrieval performance primarily stems from the MoE’s ability to capture abstract relationships across modalities through its diverse experts. Each expert specializes in particular high-level features, such as aligning the semantic content of an image with the semantic meaning of its textual description. This capability allows the model to reason beyond surface-level similarity, enabling it to uncover deeper patterns and themes within the data.
Additionally, cross-modal tasks often demand an understanding of data at multiple levels of granularity. The MoE architecture is inherently suited for hierarchical processing, with different experts focusing on varying levels of representation (e.g., object-level details in images or sentence-level semantics in text). This hierarchical modeling facilitates the construction of a structured world model, enhancing the model’s capacity for reasoning and aligning complex multimodal data effectively.
Model # Trainable Params Flickr30K (1K test set) COCO Fine-tuned (5K test set)
Image $\rightarrow$ Text Text $\rightarrow$ Image Image $\rightarrow$ Text Text $\rightarrow$ Image
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
Dual-encoder models
CLIP [37] 428M 88.0 98.7 99.4 68.7 90.6 95.2 - - - - - -
ALIGN [7] 820M 88.6 98.7 99.7 75.7 93.8 96.8 77.0 93.5 96.9 59.9 83.3 89.8
FILIP [42] 417M 89.8 99.2 99.8 75.0 93.4 96.3 78.9 94.4 97.4 61.2 84.3 90.6
Florence [44] 893M 90.9 99.1 - 76.7 93.6 - 81.8 95.2 - 63.2 85.7 -
BEIT-3 [41] 1.9B 94.9 99.9 100.0 81.5 95.6 97.8 84.8 96.5 98.3 67.2 87.7 92.8
Fusion-encoder models
UNITER [6] 303M 83.6 95.7 97.7 68.7 89.2 93.9 65.7 88.6 93.8 52.9 79.9 88.0
OSCAR [27] 345M - - - - - - 70.0 91.1 95.5 54.0 80.8 88.5
VinVL [46] 345M - - - - - - 75.4 92.9 96.2 58.8 83.5 90.3
Dual encoder + Fusion encoder
ALBEF [24] 233M 94.1 99.5 99.7 82.8 96.3 98.1 77.6 94.3 97.2 60.7 84.3 90.5
BLIP [25] 446M 97.1 100.0 100.0 86.7 97.3 98.7 82.4 95.4 97.9 65.1 86.3 91.8
BLIP-2 ViT-L [26] 474M 96.9 100.0 100.0 88.6 97.6 98.9 83.5 96.0 98.0 66.3 86.5 91.8
BLIP-2 ViT-g [26] 1.2B 97.6 100.0 100.0 89.7 98.1 98.9 85.4 97.0 98.5 68.3 87.7 92.6
LLMs + LVMs
Alt-MoE (LLaMA3-8b, DinoV2-L) 140M 97.8 100.0 100.0 97.8 100.0 100.0 87.7 99.6 99.9 89.7 99.7 99.9
Alt-MoE (Qwen2-7b, DinoV2-L) 130M 97.9 100.0 100.0 97.8 100.0 100.0 88.1 99.4 99.8 90.1 99.6 99.9
Table 1: Comparison with state-of-the-art image-text retrieval methods, finetuned on COCO and Flickr30K.
5.3 Zero-shot Audio-text Retrieval
To evaluate the model’s generalization ability on unseen data as well as its scalability across modalities and tasks, we validated Alt-MoE through zero-shot audio-to-text retrieval experiments. Specifically, we trained Alt-MoE on a pretraining dataset and validated its performance on Clotho and Audiocaps, using the same architecture as in image-text retrieval. Details of the pretraining data are provided in the supplementary materials.
As shown in the table 2, Alt-MoE demonstrates superior performance over baseline models in zero-shot retrieval tasks, showcasing its exceptional generalization capabilities. This is primarily due to two key factors: first, the model benefits from the rich intra-modal knowledge embedded in pre-trained uni-modal models, which allows it to leverage existing knowledge without the need for extensive retraining. Second, the hierarchical MoE structure significantly strengthens abstract reasoning abilities. MoE models are known to excel at capturing complex, high-level patterns by assigning specialized experts to different levels of abstraction. This structured, layered processing enables Alt-MoE to model intricate relationships between modalities, further improving its ability to generalize and make inferences based on unseen data.
MoE significantly enhances abstract reasoning by assigning specialized experts to different levels of abstraction. This allows Alt-MoE to capture complex, high-level patterns and model the intricate relationships between modalities. Such a structure improves the model’s generalization ability and inference capabilities on unseen data.
Method Clotho Audiocaps
Audio $\rightarrow$ Text Text $\rightarrow$ Audio Audio $\rightarrow$ Text Text $\rightarrow$ Audio
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
AVFIC [34] - - - 3.0 - 17.5 - - - 8.7 - 37.7
ImageBind [13] - - - 6.0 - 28.4 - - - 9.3 - 42.3
VALOR [5] - - - 8.4 - - - - - - - -
LanguageBind [47] 16.1 39.9 53.2 15.5 38.6 51.7 17.8 47.3 64.0 16.5 48.7 64.6
Alt-MoE 17.0 40.8 53.0 20.1 45.2 58.7 20.4 50.8 66.6 19.8 51.4 66.8
Table 2: Zero-shot Performance on Clotho and Audiocaps datasets.
5.4 Visual Question Answer
Visual Question Answering (VQA) is a challenging task that demands our visual language model to integrate an image with relevant questions to provide accurate responses. Our experiments are conducted on the VQAv2 dataset [14], where each image is associated with several question-answer pairs, each with distinct scores. The model processes both the image and the question concurrently. We leverage a pre-trained MOE model to synthesize features from a comprehensive set of textual and visual tokens. These tokens are encoded by the state-of-the-art Qwen2.5-7B [17] for text and the robust Dino-v2-large [35] for visual data, respectively. Following previous works [40], the aggregated output is then passed to a classifier, which predicts the answer from a pool of 3129 frequently occurring answers. We present our results based on the performance metrics obtained from the test-dev and test-std sets of the dataset. Details of the pretraining data and results Table are provided in the supplementary materials.
Alt-MoE demonstrates outstanding performance. This can be largely attributed to its dual-directional alignment mechanism and the MoE’s ability to capture dependencies across multiple levels of abstraction. Specifically, the bidirectional alignment facilitates more comprehensive understanding by allowing the model to process both image-to-question and question-to-image, ensuring that contextual and semantic nuances are preserved in both directions. Additionally, the MoE framework enhances this process by distributing the task across specialized experts, each focusing on different levels of abstraction—from object-level details to higher-level semantic reasoning. This layered reasoning capability significantly boosts the model’s performance in complex VQA tasks.
5.5 Image Classification
To clearly demonstrate the ability of multi-directional MoE to integrate existing knowledge and support hierarchical abstract reasoning, we designed an experiment that combines knowledge from various models trained on unified modalities. This experiment clearly demonstrates how MoE can facilitate complementary knowledge integration across different models. Specifically, DinoV2 [35] is a vision-only model, while CLIP-ViT [37] is a text-guided vision model. These models capture distinct types of information, offering complementary perspectives. To validate this hypothesis, we froze both DinoV2 and CLIP-ViT and extracted different representations for the same image. These representations were then aligned and fused using a MoE, followed by a classification head. As shown in the Table 3, the performance of the jointly trained models exceeds that of either model alone, confirming the complementary nature of their representations. Moreover, the vision-only model outperforms the text-guided vision model, suggesting that the latter may struggle to capture fine-grained details effectively.
Model Accuracy (%) Precision (%) Recall (%) F1 Score (%)
CLIP-ViT [37] 82.1 82.4 82.0 82.0
DinoV2 [35] 83.2 83.5 83.3 83.1
Alt-MoE 86.6 86.9 86.6 86.5
Table 3: Comparison of Classification Performance on the ImageNet-1K Dataset.
5.6 Validation of Alignment Objectives
Model Pred CL Image $\rightarrow$ Text Text $\rightarrow$ Image
R@1 R@5 R@10 R@1 R@5 R@10
Alt-MoE $\times$ 44.2 86.2 95.6 2.6 12.6 25.1
Alt-MoE 66.1 91.9 96.4 66.2 89.6 99.0
Table 4: Performance comparison of prediction alignment and joint training objectives: Image-Text retrieval with Alt-MoE (Qwen-7b, DinoV2-L) on COCO fine-tuning. CL denotes Contrastive Learning. Pred denotes prediction.
Alt-MoE utilizes the MoE router to automatically select different experts, optimizing for both objectives. This approach helps decouple modality-specific information from shared information, potentially improving the balance between capturing unique modal features and cross-modal relationships. To validate this conclusion, we applied the predicted representations to retrieval tasks. The results in Table 4 indicate that training solely on the prediction alignment objective yields significantly lower performance compared to jointly training on both prediction and contrastive learning objectives. By jointly optimizing for prediction and contrastive learning, the model maximizes the mutual information, while minimizing the conditional entropy. This ensures that Alt-MoE capture the most relevant and informative features, leading to better alignment and retrieval performance. The mutual reinforcement of the two tasks helps in effectively balancing the trade-off between capturing modality-specific features and learning cross-modal relationships.
5.7 Ablation Study
Model MLP MoE ALT Image $\rightarrow$ Text Text $\rightarrow$ Image
R@1 R@5 R@10 R@1 R@5 R@10
Alt-MoE $\times$ 74.4 86.0 92.2 82.3 89.5 92.6
Alt-MoE $\times$ $\times$ 68.2 68.7 81.1 74.2 88.7 92.4
Alt-MoE $\times$ 88.1 99.4 99.8 90.1 99.6 99.9
Table 5: Ablation Study of Image-Text Retrieval with the Alt-MoE contrastive learning representations (Qwen-7b, DinoV2-L) on COCO Fine-Tuning.
To validate its effectiveness, we conducted ablation experiments by replacing MoE with MLP and comparing the results. Additionally, we examined the impact of alternating optimization versus non-alternating optimization. As shown in Table 5, indicate that replacing MoE with MLP leads to a significant drop in performance, demonstrating the critical role of MoE. Furthermore, the comparison between alternating optimization and non-alternating optimization reveals that alternating optimization contributes to better model performance.
6 Conclusion
In this study, we introduced Alt-MoE, a novel modality-agnostic training strategy and architecture designed for multi-modal learning. Alt-MoE effectively integrates high-performance uni-modal models using bidirectional alignment module, facilitating the alignment of modality pairs in multiple directions and enabling generalization to new tasks and modalities. To validate the modality and task scalability of Alt-MoE, we conducted extensive experiments focused on alignment performance. The experimental results demonstrate that Alt-MoE can easily generalize to new modalities, tasks, and datasets while maintaining the same training strategy and architecture.